# FindingDory: A Benchmark to Evaluate Memory in Embodied Agents

### v1

@@ 6 Conclusion
We introduced FindingDory, a benchmark for evaluating long-horizon memory in embodied agents within a highly photorealistic indoor simulation. Our findings reveal limitations in VLM context scaling, particularly in handling long observation sequences required for memory-intensive tasks. FindingDory provides informative metrics that disentangle different aspects of memory reasoning, allows for procedural task generation of increasing difficulty, and fine-grained, efficiency-focused failure analysis across high-level semantic retrieval and low-level spatial reasoning. Its extensibility allows for longer temporal dependencies, supporting research on context scaling and memory efficiency in multimodal models. In our results, we observe that the supervised fine-tuning baseline performance improves consistently when trained with longer videos which suggests future work should explore finetuning VLMs with effective memory compression techniques for enhanced spatio-temporal reasoning.
Our experiments also highlight challenges with hierarchical policies that integrate long-context VLMs with low-level control. While high-level policies specify abstract goals—such as referencing a past frame in the interaction history—low-level navigation policies often struggle to reach the intended object precisely, exposing a gap between goal specification and reachability. Addressing this limitation requires better integration between high-level reasoning and low-level execution, either by improving hierarchical policies to leverage VLM guidance more effectively or by training end-to-end models that jointly learn memory retrieval and control.
Beyond evaluation, FindingDory can serve as a training platform for memory-dependent tasks, enabling long-context-aware models and enabling dataset extraction for supervised learning tasks such as VQA over long videos. By bridging the gap between VLM advancements and real-world memory demands, FindingDory facilitates systematic study of memory-augmented policy learning and long-term reasoning, driving progress in multimodal embodied AI.
@@ B.1 FindingDory Episode Creation
Object-Receptacle Pairing. The first step in the episode creation involves generating valid spawning positions for various object-receptacle pairs by sampling entities from the respective split in consideration (train/val). The objective is to identify candidate objects that can be picked up from a start receptacle and placed stably on a goal receptacle. We use the pipeline proposed in [48] to procedurally generate various object placements on receptacle surfaces by running physics checks for stable placements. To make the generation procedure easier, we assume that each object is picked and placed only once during an episode. Additionally, we constrain the start receptacle category to be distinct from the goal receptacle category. We also ensure that for each candidate object category, a corresponding non-interacted object of the same category is spawned to serve as distractor entities during the interaction phase.  Figures 5(b) and 5(c) provide a distribution of distractor objects and receptacles in the val episodes. The number of sampled object-receptacle pairings is equal to the number of rearrangements in the particular episode and can be configured through an argument. In essence, this controls the distribution of video frame lengths in the FindingDory tasks(see  Figure 5(a)).
@@
Task Creation. Once we have the final filtered set of episodes, we populate task instructions for each episode using yaml-based instruction templates as used in [38]. We employ the instruction templates to procedurally populate various placeholders such as object names, attributes and appearance orders that create specific task instructions grounded in the entities (objects and receptacles) present in the episode in an automated manner. We present the diverse instructions templates we use for each task category in Table LABEL:tab:all_tasks. In total, we create $82174/5876$ unique task instructions within the train/val split respectively.
@@ B.2 Experience Collection Phase
The above OracleAgent routine enables us to procedurally create videos of the agent rearranging objects in the environment. Additionally, as the agent navigates the environment between the pick and place substeps, it collects additional frames detailing the scene environment which can be useful (or serve as distractor frames) when attempting to solve FindingDory tasks. We present an example experience collection trajectory and associated “magic” pick-place sequence in Figures 12 and 13.
@@ B.3 Evaluations
In this section, we define the primary evaluation metrics employed to benchmark the performance of various baselines on the FindingDory tasks. Following [6], we focus on metrics that systematically quantify both the success and efficiency of the agent in solving specified instructions. Success focuses on accurately selecting frames from the interaction video that lead to task completion, while efficiency evaluates if the agent selects and navigates to the optimal frame that solves the task using the fewest low-level control actions. Notably, optimizing efficiency poses a significant challenge for current methods [29, 47].
@@
Success-weighted-by-Path-Length. This metric evaluates task completion efficiency, computed per task episode as follows [6]:
@@
gemini-2.0-flash Task Prompt You are an intelligent question answering agent. In the prompt video preceeding the text, you will be shown a long stream of images that have been collected by a robot while navigating around a home and asked to answer a question about the space to assist a user in completing tasks in this home. In the video stream, the robot is picking up objects and rearranging them to different arbitrary locations throughout the house.
The questions asked by user will require being able to look at the full set of images collected by robot to be able to provide the answer. Your task is to identify the exact point in the video (by timestamp) where the user should move to best accomplish their goal. You can do this by outputting the exact timestep in the video where you are most confident that the object or place that they should move to, was viewed closely.
The images in the video also have the time of day information in top left. You need to use the time of day information for tasks that require to revisit objects or receptacles at a specific time of day.
For some tasks that require revisiting multiple receptacles or objects, you should output multiple frame indices corresponding to the correct order of revisitation (only if specified). For all user goals, identify the minimum number of objects or receptacles (targets) that you need to revisit to complete the user’s goal. For each revisitation, you should select one corresponding timestep in the video that you will move to. You need to ensure that you are not revisiting targets that are not relevant to the user’s goal.
Generate timestamps in the MM:SS format where the first two digits represent minutes and the last two digits represent seconds. To solve the tasks effectively, you should try to summarize the video by listing out the objects that were picked and placed, the receptacles from which each object was picked up and the receptacle where it was dropped.
The user’s goal is to: {goal}. Now look at the full video containing all the images. Give your final response with NUM_TARGETS_TO_REVISIT: ||<num_target>||. TIMESTAMP_INDEX: ||<timestamp_1>,…,<timestamp_num_target>||. You can include multiple timestamps if the task requires visiting more than one target (one timestamp for each target). Do not add any other text.
Figure 6: System Prompt for gemini-2.0-flash model
@@ C.1 An overview of approaches tackling memory in Long VQA and embodied tasks
Many recent works equip large language models with memory-like capabilities through task-specific knowledge bases or semantic maps for navigation planning [25, 8]. MobilityVLA [45] leverages long-context VLMs to process past frames, enabling goal-frame selection via Gemini, while topological mapping guides navigation without explicit pathfinding. ReMEmbR [3] and Embodied RAG [43] introduce nonparametric memory trees to store and retrieve past experiences for planning. For video question answering, Zhang et al. [49], Kahatapitiya et al. [13] propose a streaming-based approach that accumulates frame-wise captions for later querying. However, this method constrains memory to the expressivity of vision-language or captioning models, potentially omitting critical details. We include the hierarchical approach involving long-context VLMs and the approach utilizing textual memory traces as baselines in our evaluations on the benchmark tasks.
@@ D.1 Gemini Agent
We experiment with the gemini-2.0-flash model which can process upto 1 million tokens in the context window. We strictly follow the outlined Google Gemini API developer guidelines [1] to structure the prompt and perform API calls over the long videos in the FindingDory tasks. We do not use a “structured output” format such as json in the output responses as we empirically observed it failed to produce coherent outputs when supplied with long videos in the context. Since the gemini-2.0-flash model expects the input frames to be sampled at 1 FPS, we prompt the model to directly generate specific timestamps in the recommended MM:SS format to localize the frame corresponding to the target entity. We provide the detailed prompt that we use for evaluations in Fig. 6.
D.2 Qwen and Gemma Agent
@@
We run the Qwen2.5-VL-7B [5] and Gemma3-12B [40] locally for evaluations on the FindingDory tasks. The Qwen model uses a maximum of $768$ frames sampled uniformly across the input sequence while Gemma does not impose a maximum frame limit. For evaluation on 768 frames, we use two A40 GPUs with 48GB of memory. For experiments with shorter context lengths and higher subsampling (24–384 frames), a single A40 GPU suffices. We provide the full task prompt we use for inference with Qwen and Gemma in Fig. 7 and utilize structured JSON outputs.
@@
Qwen/Qwen2.5-VL-7B-Instruct and google/gemma-3-12b-it Task Prompt You are an expert and intelligent question answering agent. You will be shown a video that was collected by a robot yesterday while navigating around a house and picking and placing objects. Your job is to help the robot complete a task today by looking at the video and finding the frame indices that the robot should move to. Make sure your response is all within the JSON.
Note:
- The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
Output Format (should be in JSON format):
@@
Where:
- Chain of Thought: A detailed explanation of the robot’s thought process in determining the frame indices.
- Frame Indices: The frame indices in the video stream showing the object or place that the robot should move to.
- When deciding which frame indices to choose, make sure you choose the frame indices that are closest to the object/place.
- If task requires the agent to go to multiple places, output one frame index per object/place. (Do not use ellipsis)
The robot’s goal is: {goal}
Figure 7: System Prompt for Qwen2.5-VL and Gemma3 model
D.3 GPT Agent
We use the official OpenAI GPT-4o [12] API to run evaluations on the FindingDory tasks with video frames subsampled at $96$ frames to optimize API costs. For running complete evaluation on the entire validation task suite, we incurred a total cost of $\approx 400$ USD. We provide the full system prompt used for GPT-4o evaluations in Fig. 8.
gpt-4o-2024-07-18 Task Prompt ** OBJECTIVE **
You are an expert and intelligent question answering agent. You will be shown a video that was collected by a robot yesterday while navigating around a house and picking and placing objects.
Your job is to help the robot complete a task today by looking at the video and finding the frame indices that the robot should move to, to complete the task.
** OUTPUT FORMAT **
@@ D.3 GPT Agent
Where:
- Chain of Thought: A detailed explanation of the robot’s thought process in determining the frame indices.
- Frame Indices: The frame indices in the video stream showing the object or place that the robot should move to.
- When deciding which frame indices to choose, make sure you choose the frame indices that are closest to the object/place.
- If task requires the agent to go to multiple places, output one frame index per object/place.
** NOTES **
- The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
** TASK **
The robot’s goal is: {goal}
Figure 8: System Prompt for gpt-4o model

### v2

@@ 6 Conclusion
We introduced FindingDory, a benchmark for evaluating long-horizon memory in embodied agents within a highly photorealistic indoor simulation. Our findings reveal limitations in VLM context scaling, particularly in handling long observation sequences required for memory-intensive tasks. FindingDory provides informative metrics that disentangle different aspects of memory reasoning, allows for procedural task generation, and efficiency-focused failure analysis beyond simple frame selection accuracy. Its extensibility allows for longer temporal dependencies, supporting research on context scaling and memory efficiency in multimodal models. Our experiments also highlight challenges with hierarchical policies motivating the need for better architectures that closely integrate long-context VLM capabilities with generalist embodied policies.
@@ B.1 FindingDory Episode Creation
Object-Receptacle Pairing. The first step in the episode creation involves generating valid spawning positions for various object-receptacle pairs by sampling entities from the respective split in consideration (train/val). The objective is to identify candidate objects that can be picked up from a start receptacle and placed stably on a goal receptacle. We use the pipeline proposed in Yenamandra et al. (2023) to procedurally generate various object placements on receptacle surfaces by running physics checks for stable placements. To make the generation procedure easier, we assume that each object is picked and placed only once during an episode. Additionally, we constrain the start receptacle category to be distinct from the goal receptacle category. We also ensure that for each candidate object category, a corresponding non-interacted object of the same category is spawned to serve as distractor entities during the interaction phase.  Figures 5(b) and 5(c) provide a distribution of distractor objects and receptacles in the val episodes. The number of sampled object-receptacle pairings is equal to the number of rearrangements in the particular episode and can be configured through an argument. In essence, this controls the distribution of video frame lengths in the FindingDory tasks(see  Figure 5(a)).
@@
Task Creation. Once we have the final filtered set of episodes, we populate task instructions for each episode using yaml-based instruction templates as used in  Szot et al. (2024). We employ the instruction templates to procedurally populate various placeholders such as object names, attributes and appearance orders that create specific task instructions grounded in the entities (objects and receptacles) present in the episode in an automated manner. We present the diverse instructions templates we use for each task category in Table LABEL:tab:all_tasks. In total, we create $82174/5876$ unique task instructions within the train/val split respectively.
@@ B.2 Experience Collection Phase
The above OracleAgent routine enables us to procedurally create videos of the agent rearranging objects in the environment. Additionally, as the agent navigates between the pick and place location, it collects additional frames detailing the scene environment which can be useful (or serve as distractor frames) when attempting to solve FindingDory tasks. We present an example experience collection trajectory and associated “magic” pick-place sequence in Figures 16 and 17.
@@ B.3 Evaluations
In this section, we define the primary evaluation metrics employed to benchmark the performance of various baselines on the FindingDory tasks. Following Batra et al. (2020), we focus on metrics that systematically quantify both the success and efficiency of the agent in solving specified instructions. Success focuses on accurately selecting frames from the interaction video that lead to task completion, while efficiency evaluates if the agent selects and navigates to the optimal frame that solves the task using the fewest low-level control actions. Notably, optimizing efficiency poses a significant challenge for current methods (Ramakrishnan et al., 2024; Yang et al., 2024).
@@
Success-weighted-by-Path-Length. This metric evaluates task completion efficiency, computed per task episode as follows (Batra et al., 2020):
@@
B.4 Subsampled Videos Solvability Analysis
Figure 6: Task Solvability Ratio at the $96$ frame subsampling level.
In this section, we analyze the solvability of the tasks proposed in FindingDory suite when all videos are subsampled to enable full-finetuning of VLMs (see section D.5 for details). Specifically, we conduct a solvability analysis at the $96$-frame video subsampling level. Full finetuning of the VLM on these videos enables highest task performance of $\approx 52\%$ as shown in  fig. 4(a).
From  fig. 6 we observe that after subsampling the videos to $96$ frames, there is a $\approx 1\%$ drop in solvability across the various low-level task categories. The largest drop is observed in the Conditional Interaction tasks where a specific object or receptacle is to be revisited based on the original interaction sequence. We want to highlight that this solvability calculation is solely based on whether a solution frame (from multiple possible frames in the full length video) is left over in the subsampled video. Such frames are treated as a solution based on whether they satisfy the PDDL goal conditions (see  section 3.2) corresponding to the task. But such frames are extremely difficult to select for goal completion for a VLM since the subsampling routine causes a large drop in the relevant context that is needed to make sense of the solution frame. For instance, consider the task “Navigate to the object which you interacted with which is the farthest from your current location”: it is possible that a frame with the target object is left over in the subsampled video. But selecting this solution frame is nearly impossible since the subsampling to $96$ frames eliminates all the spatiotemporal context that is necessary to solve this task.
From  fig. 6 we also observe that $3.85\%$ of the tasks are Originally Unsolvable which correspond to tasks focused on distractor entities such as “Navigate to a soft toy that you did not rearrange yesterday” or “Navigate to a room that you did not visit yesterday”.
B.5 Chain-of-Thought Trace Generation
Figure 7: Textual Memory Trace corresponding to agent interaction during Experience Collection. The memory trace is subsampled to ensure alignment with the 48-frame video.
In this section, we outline the pipeline that is employed to generate detailed chain-of-thought traces for the FindingDory training data split. We begin by generating textual memory traces that detail the interactions performed by the robot during the Experience Collection phase (see Section B.2). For this, we track all the individual actions that are executed by the robot along with necessary information such as time of day, object-recetpacle info, room names that are necessary for successful task completion. For the co-training experiments for cross-benchmark analysis, we subsample all videos to $48$ frames and accordingly modify the textual memory traces to ensure alignment between the memory trace and subsampled video frames. A sample original and subsampled textual memory trace is shown in Figure 7.
To generate faithful chain-of-thought traces for all samples in the training split, we leverage a zero-shot, text-based LLM to generate a rationale for each task based on the textual memory history and corresponding keyframe solution. We hypothesize that the zero-shot LLM would be able to generate more consistent traces if it is given access to the solution keyframes and prompted to generate a rationale using the solution. This is in contrast to using the LLM to predict the solution keyframes based on the textual memory trace which is a more difficult task. We use the open-source DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) for this purpose and leverage the vLLM library (Kwon et al., 2023) for accelerated inference. We use the prompt structure as shown in  Figure 8. After the initial chain-of-thought trace generation, we observed that $\approx 8\%$ of the generated traces contained references to the solution keyframe list provided in the prompt. To fix these erroneous traces, we use the DeepSeek-R1-Distill-Qwen-32B model to rephrase the generated traces to remove any references to the solution keyframe list following the prompt shown in Figure 9. After the initial trace generation and rephrasing steps, we are left with $7.8\%$ traces in which the solution keyframes list predicted by the LLM contains a keyframe which does not belong to the actual solution keyframes list. All these samples are discarded which leads to a total of $72973$ training samples with labeled chain-of-thought traces.
DeepSeek-R1-Distill-Qwen-32B CoT Generation Prompt You are an intelligent agent assisting a robot in completing specific goals by analyzing interactions recorded during its previous navigation around a house. The interaction history captures the robot rearranging objects (pick and place). The robot is given a specific task that needs to be solved by selecting a specific frame index between 0-48. Each task is solved by selecting the correct frame from the history. Your task is to help the robot generate a detailed reasoning chain of thought to solve the task. To aid in generating the reasoning trace, you will be given the full list of oracle keyframes that will solve the task. Choosing any frame from this list solves the task. If the oracle list contains only [-1] element, then the task is just unsolvable. The reasoning_trace should terminate with a list of predicted solution keyframes. Use the oracle solution to verify your solution_ keyframes predicted list. You need to solve the following task: {task_instruction}. The oracle solution is: {oracle_solution}. The detailed interactions of the robot are: {textual_memory_trace} Explanation of oracle_solution list format: - Single list with multiple sublists. Each sublist can contain multiple frame indices. - If only one sublist, then the task is a single-goal task - If multiple sublists, then the tasks has multiple subgoals and each sublist has corresponds to the solution frames of respective subgoal Please generate the detailed reasoning trace and make sure you refer to the various frame indices and robot actions in the interaction history. Guidelines for reasoning_trace generation: - Do not include explicit references to the oracle keyframes. - Your answer should always end with the list of frames that solve the task (use [-1] if unsolvable) based on the reasoning trace. - Ensure the predicted solution_keyframes follows the format of the oracle_solution list. - Make sure that each value in the final list of solution_keyframes you generated exists in the oracle_solution list that is provided. Please follow the output format: Output Format (should be in JSON format):
  "reasoning_trace": str.
  "solution_keyframes": str.
Figure 8: Prompt used for CoT trace generation with DeepSeek-R1-Distill-Qwen-32B.
DeepSeek-R1-Distill-Qwen-32B CoT Trace Rephrasing You are given a task instruction and a reasoning trace (chain of thought) that was generated to solve the task. The reasoning trace currently contains explicit reference to “oracle solution list” or similar oracle-based information that should not be visible to a model trying to solve the task independently. Your job is to rephrase the reasoning trace to remove any explicit references to oracle solutions or oracle keyframes, while preserving the core reasoning logic and thought process. The rephrased reasoning should read as if it was generated by a model thinking through the problem step by step without access to ground truth answers. You should keep references to all other details such as individual frame indices or segments intact as they are necessary for reasoning. Task Instruction: {task_instruction} Original Chain of Thought: {existing_chain_of_thought} Please follow the output format: Output Format (should be in JSON format): “‘
  "reasoning_trace": str.
Figure 9: Prompt for CoT Trace Rephrasing with DeepSeek-R1-Distill-Qwen-32B.
B.6 Sanity Checks
We investigate how a VLM would perform at short-horizon visual comprehension on images from the same episodes used for long-horizon evaluation in our benchmark. This tests if poor performance of the evaluated models on our long-horizon reasoning tasks is due to a domain or sim2real gap. We assess this by sampling frames around object interactions, asking models to list out detected objects, and check accuracy compared to the ground truth category via string matching – a conservative metric that misses correct predictions using synonyms or alternative phrasings. This still results in image sequences of length 50-150, but Gemini achieves an accuracy of 67% which is significantly higher than it’s average high level success rate across tasks in fig. 2. This shows that despite clear, close-up views, models struggle to reason over them when temporal sequence comprehension is required.
@@ C.1 An overview of approaches tackling memory in Long VQA and embodied tasks
Many recent works equip large language models with memory-like capabilities through task-specific knowledge bases or semantic maps for navigation planning (Min et al., 2024; Chang et al., 2023). MobilityVLA (Xu et al., 2024) leverages long-context VLMs to process past frames, enabling goal-frame selection via Gemini, while topological mapping guides navigation without explicit pathfinding. ReMEmbR (Anwar et al., 2024) and Embodied RAG (Xie et al., 2024) introduce nonparametric memory trees to store and retrieve past experiences for planning. For video question answering, Zhang et al. (2023); Kahatapitiya et al. (2024) propose a streaming-based approach that accumulates frame-wise captions for later querying. However, this method constrains memory to the expressivity of vision-language or captioning models, potentially omitting critical details. We include the hierarchical approach involving long-context VLMs and the approach utilizing textual memory traces as baselines in our evaluations on the benchmark tasks.
@@ D.1 Gemini Agent
gemini-2.0-flash Task Prompt You are an intelligent question answering agent. In the prompt video preceeding the text, you will be shown a long stream of images that have been collected by a robot while navigating around a home and asked to answer a question about the space to assist a user in completing tasks in this home. In the video stream, the robot is picking up objects and rearranging them to different arbitrary locations throughout the house. The questions asked by user will require being able to look at the full set of images collected by robot to be able to provide the answer. Your task is to identify the exact point in the video (by timestamp) where the user should move to best accomplish their goal. You can do this by outputting the exact timestep in the video where you are most confident that the object or place that they should move to, was viewed closely. The images in the video also have the time of day information in top left. You need to use the time of day information for tasks that require to revisit objects or receptacles at a specific time of day. For some tasks that require revisiting multiple receptacles or objects, you should output multiple frame indices corresponding to the correct order of revisitation (only if specified). For all user goals, identify the minimum number of objects or receptacles (targets) that you need to revisit to complete the user’s goal. For each revisitation, you should select one corresponding timestep in the video that you will move to. You need to ensure that you are not revisiting targets that are not relevant to the user’s goal. Generate timestamps in the MM:SS format where the first two digits represent minutes and the last two digits represent seconds. To solve the tasks effectively, you should try to summarize the video by listing out the objects that were picked and placed, the receptacles from which each object was picked up and the receptacle where it was dropped. The user’s goal is to: {goal}. Now look at the full video containing all the images. Give your final response with NUM_TARGETS_TO_REVISIT: ||<num_target>$|$. TIMESTAMP_INDEX: ||<timestamp_1>,…,<timestamp_num_target>$|$. You can include multiple timestamps if the task requires visiting more than one target (one timestamp for each target). Do not add any other text.
Figure 10: Prompt for gemini-2.0-flash model
@@
We experiment with the gemini-2.0-flash model which can process upto 1 million tokens in the context window. We strictly follow the outlined Google Gemini API developer guidelines (gem, ) to structure the prompt and perform API calls over the long videos in the FindingDory tasks. We do not use a “structured output” format such as json in the output responses as we empirically observed it failed to produce coherent outputs when supplied with long videos in the context. Since the gemini-2.0-flash model expects the input frames to be sampled at 1 FPS, we prompt the model to directly generate specific timestamps in the recommended MM:SS format to localize the frame corresponding to the target entity. We provide the detailed evaluations prompt in Fig. 10.
D.2 Qwen, Gemma and GLM Agent
@@
We run the Qwen2.5-VL-7B (Bai et al., 2025), Gemma3-12B (Team et al., 2025) and GLM-4.1V-Thinking (Hong et al., 2025) locally for evaluations on the FindingDory tasks. The Qwen model uses a maximum of $768$ frames sampled uniformly across the input sequence while Gemma and GLM does not impose a maximum frame limit. For evaluation on 768 frames, we use two A40 GPUs with 48GB of memory. For experiments with shorter context lengths and higher subsampling (24–384 frames), a single A40 GPU suffices. We provide the full task prompt we use for inference with the three agents in Fig. 11 and utilize structured JSON outputs.
@@
Qwen/Qwen2.5-VL-7B-Instruct, google/gemma-3-12b-it and zai-org/GLM-4.1V-9B-Thinking Task Prompt You are an expert and intelligent question answering agent. You will be shown a video that was collected by a robot yesterday while navigating around a house and picking and placing objects. Your job is to help the robot complete a task today by looking at the video and finding the frame indices that the robot should move to. Make sure your response is all within the JSON. Note: - The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up. Output Format (should be in JSON format):
@@
Where: - Chain of Thought: A detailed explanation of the thought process in determining the frame indices. - Frame Indices: The frame indices in the video stream showing the object or place that the robot should move to. - When deciding which frame indices to choose, make sure you choose the frame indices that are closest to the object/place. - If task requires the agent to go to multiple places, output one frame index per object/place. (Do not use ellipsis) The robot’s goal is: {goal}
Figure 11: Prompt for Qwen2.5-VL and Gemma3 model
gpt-4o-2024-07-18 Task Prompt ** OBJECTIVE ** You are an expert and intelligent question answering agent. You will be shown a video that was collected by a robot yesterday while navigating around a house and picking and placing objects. Your job is to help the robot complete a task today by looking at the video and finding the frame indices that the robot should move to, to complete the task. ** OUTPUT FORMAT **
@@
Where: - Chain of Thought: A detailed explanation of the thought process for determining the frame indices. - Frame Indices: The frame indices in the video stream showing the object or place that the robot should move to. - When deciding which frame indices to choose, make sure you choose the frame indices that are closest to the object/place. - If task requires the agent to go to multiple places, output one frame index per object/place. ** NOTES ** - The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up. ** TASK ** The robot’s goal is: {goal}
Figure 12: Prompt for gpt-4o model
D.3 GPT Agent
We use the official OpenAI GPT-4o (Hurst et al., 2024) API to run evaluations on the FindingDory tasks with video frames subsampled at $96$ frames to optimize API costs. For running complete evaluation on the entire validation task suite, we incurred a total cost of $\approx 400$ USD. We provide the full system prompt used for GPT-4o evaluations in Fig. 12.